1) a. Find the probability of rolling a total of at least 11 with two dice.

b. Find the probability of choosing, without replacement, a spade and then a heart out of a standard deck of cards.
Answer #1

Answer)

A)

Probability is given by favorable / total

Total = 6*6 = 36

Favorable = sum of at least 11

(5, 6), (6, 6), (6, 5) = 3

Probability = 3/36 = 0.08333333333

B)

Total = 52

Favorable = spade = 13

P(spade) = 13/52

P(heart) = 13/51 (as are not replacing the previous one, so we have 13 hearts and total 51 cards)

P(spade then heart) = (13/52)*(13/51) = 0.06372549019